OpenStudy (anonymous):

Solve The Equation 7 4/7 = n + 4 3/7

OpenStudy (anonymous):

can you draw it?

OpenStudy (anonymous):

http://openstudy.com/study#/updates/5046e900e4b0215a1ea2ccd5 Again not sure if you mean 4*3/7 or 4 + 3/7

OpenStudy (anonymous):

\[4\frac{ 3 }{ 7 } = n + 4 \frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

oops the first fraction i made a mistake but i hope you get it